# Soft symmetries of topological orders

### v3

@@ IV.1 Quantum double model
The local Hilbert space on each edge has $|{G_{\mathrm{sf}}}|=2^{7}$ dimensions, whose basis states {|$\{\ket{g}\}$ are labeled by group elements $g\in{G_{\mathrm{sf}}}$. The Hamiltonian is given by
@@
$A_{v}=\frac{1}{|{G_{\mathrm{sf}}}|}\sum_{g\in{G_{\mathrm{sf}}}}\overrightarrow{X}_{N(v)}^{g}\overrightarrow{X}_{E(v)}^{g}\overleftarrow{X}_{W(v)}^{g^{-1}}\overleftarrow{X}_{S(v)}^{g^{-1}},\quad B_{f}=\delta_{g_{01}g_{13}g^{-1}_{02}g^{-1}_{23},0}~{}.$ (39)
@@
since $\omega_{pq}(0123)=\pi(n_{p}(01)n_{q}(13)+n_{p}(02)n_{q}(23))$. [^5] See e.g., [56] for the details of cup product on square and hypercubic lattices. Therefore, this operator corresponds to the action of the gauged SPT operator $e^{i\int\omega_{pq}}$ acting on the flat $G_{\text{sf}}$ gauge fields on a 2d space. $\mathcal{R}_{\mathbb{Z}_{2}}$ remains unchanged on a closed surface if we use a different cocycle representative for $\omega$. In particular, since we have $[\omega_{pq}]=[\omega_{\text{sf}}]$ in cohomology $H^{2}(B{G_{\mathrm{sf}}},U(1))$, the above $\mathcal{R}_{{\mathbb{Z}}_{2}}$ can also be expressed as $\mathcal{R}_{{\mathbb{Z}}_{2}}=\prod_{f=(0123)}e^{i\omega_{\text{sf}}(0123)}$, which corresponds to an alternative way of decomposing the finite depth circuit $\mathcal{R}_{\mathbb{Z}_{2}}$ into local unitaries.
@@
IV.1.1 Trivial permutation action on anyons
@@
where $C$ labels the magnetic flux, and $R$ labels the electric charge attached to it. One can check that the symmetry operator $\mathcal{R}_{{\mathbb{Z}}_{2}}$ preserves the above closed ribbons $K_{\sigma}^{RC}$. For instance, let us take a conjugacy class $C=[p]$ where $p,q\in{G_{\mathrm{sf}}}$ is introduced in (32). In that case, the ${\mathbb{Z}}_{2}$ symmetry acts on $K^{R,[p]}_{\sigma}$ by attaching the string of Pauli $Z^{(2)}$ operators parallel to the ribbon $\sigma$,
@@
IV.1.2 Symmetry action on anyon junctions
@@
where the product is over the junctions of magnetic fluxes. We note that since $[\omega_{pq}]=[\omega_{\text{sf}}]$ in cohomology $H^{2}(B{G_{\mathrm{sf}}},U(1))$, the above action on the closed ribbon network can be expressed as
@@
IV.1.3 Emergent symmetry and faithful action on genus $g\geq 2$ surfaces
@@
Meanwhile, $\mathcal{R}_{\mathbb{Z}_{2}}$ defines a non-trivial operator on the ground state subspace of a genus two surface. For instance, let us consider the genus two surface where each pair of cycles carries the ${G_{\mathrm{sf}}}$ holonomy $(p,q)$, $(r_{1},r_{2})$. Since $pqp^{-1}q^{-1}=r_{1}r_{2}r_{1}^{-1}r_{2}^{-1}=x$, this defines a nontrivial ${G_{\mathrm{sf}}}$ gauge field. See Fig. 8. The operator $\mathcal{R}_{{\mathbb{Z}}_{2}}$ then evaluates to a $(-1)$ phase on the state with this ${G_{\mathrm{sf}}}$ gauge field configuration. In general, the symmetry action on the Hilbert space is a diagonal matrix with elements $1$ or $-1$ which are partition functions of the (1+1)D $\omega_{\text{sf}}$ SPT with the given ${G_{\mathrm{sf}}}$ background.
@@ V Distinct gapped boundaries with identical condensed particles
Let us comment about the characterization of the Lagrangian algebra in terms of modularity. It is known that the Lagrangian algebra object $\mathcal{L}$ satisfies the modularity condition; suppose we have $\mathcal{L}=\bigoplus_{a}Z_{0a}a$ with simple anyons $a$ and non-negative coefficients $\{Z_{0a}\}$, then the coefficients of the Lagrangian algebra object satisfy $SZ=Z,TZ=Z$ with modular $S,T$ matrices [60]. These equations are derived by putting the gapped boundary on the torus, and studying the modular invariance of the corresponding boundary state [61]. However, since the above conditions do not rely on the multiplication morphism $\mu$, this does not give a faithful characterization of the Lagrangian algebra; it does not distinguish the gapped boundaries $({G_{\mathrm{sf}}},0)$ and $({G_{\mathrm{sf}}},\omega_{\text{sf}})$. Even before that, it is known that the above modularity condition is not a sufficient condition for the Lagrangian algebra object $\mathcal{L}$ [62]. One can extend the above modularity conditions to the higher-genus case by considering the gapped boundary at generic Riemann surface, and studying the invariance of the corresponding boundary state under mapping class groups [61] [^7] Here, the boundary state is constructed by considering the TQFT on $\Sigma_{g}\times I$ where $\Sigma_{g}$ is the genus $g$ surface and $I$ is the interval, and then one end of the interval is filled with the gapped boundary. This defines a TQFT state on $\Sigma_{g}$ on the other end, which is the boundary state. Due to the topological nature of the gapped boundary, the boundary state has to be invariant under the mapping class group actions of $\Sigma_{g}$. However, such family of conditions within the genus $<g$ still does not fully characterize the Lagrangian algebra. For instance, there are (1+1)D SPT phases with a certain finite group $G^{(g)}_{\text{sf}}$ which cannot be detected on the genus $<g$ surface, but can be detected on genus $g$ (see Appendix B). This leads to a pair of gapped boundaries of the $G_{\text{sf}}^{(g)}$ gauge theory $(G_{\text{sf}}^{(g)},0),(G_{\text{sf}}^{(g)},\omega)$ that cannot be distinguished from the boundary state with genus $<g$.
@@ V.1 Application: Oblique SSB phase of $\mathrm{Rep}(G_{\mathrm{sf}})$ symmetry in (1+1)D
Nevertheless, these two phases with fully broken $\text{Rep}(G_{\text{sf}})$ are distinct phases. In particular, if one SSB phase is continuously connected to the other by a deformation preserving $\text{Rep}(G_{\text{sf}})$ symmetry, it has to go through a phase transition. This can be most easily seen by gauging the $\text{Rep}(G_{\text{sf}})$ symmetry [68], then the path connecting the SSB phases becomes the path from a trivial to nontrivial SPT phase preserving the dual $G_{\text{sf}}$ symmetry. This has to experience a phase transition. [^8] We thank Nat Tantivasadakarn for helpful discussions on this point.
@@ VI Higher dimensions: $Q_{8}$ gauge theory in (3+1)D
Since this slant product is trivial in cohomology $H^{2}(B{\mathbb{Z}}_{4},U(1))$, this $\nu=1$ phase does not induce permutation of the magnetic flux $[\mathbf{i}$]. However, the above slant product leads to a subtle algebraic mixture among SPT defects, the magnetic fluxes and Wilson lines [50]. Suppose we have the junction of four SPT defects labeled by $\nu=1$ into $\nu=4$. If the magnetic surface operator labeled by the conjugacy class $[\mathbf{i}]$ intersects with this junction, the intersection has to emit a ${\mathbb{Z}}_{4}$ electric Wilson line $\exp(\frac{2\pi i}{4}\int a)$ carrying the charge of ${\mathbb{Z}}_{4}=Z(\mathbf{i})$. Since only the 2d irreducible representation of $Q_{8}$ can carry the ${\mathbb{Z}}_{4}$ charge under $Z(\mathbf{i})$, one can say that this junction emits the 2d Wilson line of $Q_{8}$. This is a non-invertible categorical symmetry that involves the junction of 0-form SPT defects, magnetic fluxes and the 2d non-invertible Wilson line. See Fig. 9 for illustrations. The whole discussion about the $\nu=1$ defect above is symmetric under permuting group elements among $\mathbf{i},\mathbf{j},\mathbf{k}$.
@@
More generally, when $G$ is a discrete subgroup of $SU(2)$, one can always find a topological response $\omega\in Z^{3}(BG,U(1))$ which is nontrivial in cohomology $[\omega]\neq 0$, but its slant product with any group element $g\in G$ becomes completely trivial $i_{g}\omega=0$ at the cochain level [70]. This implies that this class of finite groups has an analogue of soft symmetry, whose only nontrivial feature is an action on the point junctions of magnetic surfaces by a phase factor.
Figure 9: The junction of four $\nu=1$ defects into a $\nu=4$ defect. When the non-invertible magnetic surface operator $[\mathbf{i}]=[\mathbf{j}]=[\mathbf{k}]$ intersects with this junction, their intersection emits an electric charge carried by the 2d Wilson line of $Q_{8}$ gauge theory.
@@ VII Discussion
Finally, it remains unknown if the generic soft symmetry has the $H^{3}$ obstruction characterized by $[\mathcal{O}_{3}]=\mathcal{H}^{3}(\text{Aut}(\mathcal{C}),\mathcal{A})$. While the example of the soft symmetry provided in the main text does not have the $H^{3}$ obstruction, it remains unclear if the generic soft symmetries including the outer automorphism $\text{Out}_{\text{sf}}(G)$ carry the trivial $H^{3}$ obstruction or not. [^10] See Refs. [3, 72] for examples of the outer automorphism in finite gauge theories with $H^{3}$ obstructions. This is an interesting question since it is often believed that the symmetry that does not permute anyons leads to trivial $H^{3}$ obstruction [3, 41].
@@ Appendix A Symmetric representative of second cohomology $H^{2}(B{G_{\mathrm{sf}}},U(1))$
Note that $\omega_{\text{sf}}$ gives another representative of $H^{2}(B{G_{\mathrm{sf}}},U(1))$ with $[\omega_{\text{sf}}]=[\omega_{pq}]$. Let us explicitly construct such a function $\chi$. For convenience, a group element $g\in{G_{\mathrm{sf}}}$ is sometimes expressed as $g=g_{A}g_{B}$, where $g_{A}=p^{\alpha}q^{\beta}r_{1}^{\gamma_{1}}r_{2}^{\gamma_{2}}$ and $g_{B}=x^{\mu}y_{1}^{\nu_{1}}y_{2}^{\nu_{2}}$. We define a subset $\Gamma_{A}\subset{G_{\mathrm{sf}}}$ that consists of group elements $\Gamma_{A}=\{p^{\alpha}q^{\beta}r_{1}^{\gamma_{1}}r_{2}^{\gamma_{2}}\}$ with generic choices of $\alpha,\beta,\gamma_{1},\gamma_{2}$. Let us similarly define the subgroup $\Gamma_{B}\subset{G_{\mathrm{sf}}}$ where $\Gamma_{B}=\{x^{\mu}y_{1}^{\nu_{1}}y_{2}^{\nu_{2}}\}$.
@@
We consider the group $G^{(g)}_{\text{sf}}$ of order $2^{g^{2}+2g-1}$ with $g\geq 2$ which is the central extension ${\mathbb{Z}}_{2}^{g^{2}-1}\to G^{(g)}_{\text{sf}}\to{\mathbb{Z}}_{2}^{2g}$. The generators of ${\mathbb{Z}}_{2}^{2g}$ are labeled by $p_{j},q_{j}$ with $1\leq j\leq g$, and generators of ${\mathbb{Z}}_{2}^{g^{2}-1}$ are labeled by $x_{j}$ with $11\leq j\leq g-$ 1, and $y_{jk},z_{jk}$ with $1\leq j<k\leq g$. The central extension is characterized by
[…]
where $[p,q]=pqp^{-1}q^{-1}$ is a group commutator. Note that the above group $G^{(g)}_{\text{sf}}$ with $g=2$ corresponds to the group ${G_{\mathrm{sf}}}$ discussed in the main text. Let us consider the SPT action $\omega^{(g)}\in Z^{2}(BG^{(g)}_{\text{sf}},U(1))$ characterized by
@@
where $n_{p_{1}}(h)\in Z^{1}(BG^{(g)}_{\text{sf}},U(1))$ is the mod 2 number of $p_{1}$ that appears in the expression of $h\in G^{(g)}_{\text{sf}}$ in terms of the product of group generators. Similar for $n_{q_{1}}(h)$. This SPT phase is detected on a genus $g$ surface with holonomy $(p_{j},q_{j})$, $1\leq j\leq g$ on each pair of cycles, while cannot be detected on closed oriented surfaces of genus $<g$.

### v4

@@ IV.1 Quantum double model
[…]
@@
$A_{v}=\frac{1}{|{G_{\mathrm{sf}}}|}\sum_{g\in{G_{\mathrm{sf}}}}\overrightarrow{X}_{N(v)}^{g}\overrightarrow{X}_{E(v)}^{g}\overleftarrow{X}_{W(v)}^{g^{-1}}\overleftarrow{X}_{S(v)}^{g^{-1}},\quad B_{f}=\delta_{g_{01}g_{13}g^{-1}_{23}g^{-1}_{02},0}~.$ (39)
@@
since $\omega_{pq}(0123)=\pi(n_{p}(01)n_{q}(13)+n_{p}(02)n_{q}(23))$. [^5] See e.g., [56] for the details of cup product on square and hypercubic lattices. Therefore, this operator corresponds to the action of the gauged SPT operator $e^{i\int\omega_{pq}}$ acting on the flat $G_{\text{sf}}$ gauge fields on a 2d space. $\mathcal{R}_{\mathbb{Z}_{2}}$ remains unchanged on a closed surface if we use a different cocycle representative for $\omega$. In particular, since we have $\omega_{pq}]=[\omega_{\text{sf}}]$ in cohomology $H^{2}(B{G_{\mathrm{sf}}},U(1))$, the above $\mathcal{R}_{{\mathbb{Z}}_{2}}$ can also be expressed as $\mathcal{R}_{{\mathbb{Z}}_{2}}=\prod_{f=(0123)}e^{i\omega_{\text{sf}}(0123)}$, which corresponds to an alternative way of decomposing the finite depth circuit $\mathcal{R}_{\mathbb{Z}_{2}}$ into local unitaries.
@@
IV.2 Trivial permutation action on anyons
@@
where $C$ labels the magnetic flux, and $R$ labels the electric charge attached to it. This corresponds to the closed string of the anyon labeled by a pair $(C,R)$ as introduced in (30).
One can check that the symmetry operator $\mathcal{R}_{{\mathbb{Z}}_{2}}$ preserves the above closed ribbons $K_{\sigma}^{RC}$. For instance, let us take a conjugacy class $C=[p]$ where $p,q\in{G_{\mathrm{sf}}}$ is introduced in (32). In that case, the ${\mathbb{Z}}_{2}$ symmetry acts on $K^{R,[p]}_{\sigma}$ by attaching the string of Pauli $Z^{(2)}$ operators parallel to the ribbon $\sigma$,
@@
IV.3 Symmetry action on anyon junctions
@@
where the product is over the junctions of magnetic fluxes. We note that since $\omega_{pq}]=[\omega_{\text{sf}}]$ in cohomology $H^{2}(B{G_{\mathrm{sf}}},U(1))$, the above action on the closed ribbon network can be expressed as
@@
IV.4 Emergent symmetry and faithful action on genus $g\geq 2$ surfaces
@@
Meanwhile, $\mathcal{R}_{\mathbb{Z}_{2}}$ defines a non-trivial operator on the ground state subspace of a genus two surface. For instance, let us consider the genus two surface where each pair of cycles carries the $G_{\mathrm{sf}}$} holonomy $(p,q)$, $(r_{1},r_{2})$. Since $pqp^{-1}q^{-1}=r_{1}r_{2}r_{1}^{-1}r_{2}^{-1}=x$, this defines a nontrivial $G_{\mathrm{sf}}$} gauge field. See Fig. 8. The operator $\mathcal{R}_{{\mathbb{Z}}_{2}}$ then evaluates to a (1)(-1) phase on the state with this $G_{\mathrm{sf}}$} gauge field configuration. In general, the symmetry action on the Hilbert space is a diagonal matrix with elements $1$ or 1-1 which are partition functions of the (1+1)D $\omega_{\text{sf}}$ SPT with the given $G_{\mathrm{sf}}$} background.
@@ V Distinct gapped boundaries with identical condensed particles
Let us comment about the characterization of the Lagrangian algebra in terms of modularity. It is known that the Lagrangian algebra object $\mathcal{L}$ satisfies the modularity condition; suppose we have $\mathcal{L}=\bigoplus_{a}Z_{0a}a$ with simple anyons $a$ and non-negative coefficients $\{Z_{0a}\}$, then the coefficients of the Lagrangian algebra object satisfy $SZ=Z,TZ=Z$ with modular $S,T$ matrices [60]. These equations are derived by putting the gapped boundary on the torus, and studying the modular invariance of the corresponding boundary state [61]. However, since the above conditions do not rely on the multiplication morphism $\mu$, this does not give a faithful characterization of the Lagrangian algebra; it does not distinguish the gapped boundaries $({G_{\mathrm{sf}}},0)$ and $({G_{\mathrm{sf}}},\omega_{\text{sf}})$. Even before that, it is known that the above modularity condition is not a sufficient condition for the Lagrangian algebra object $\mathcal{L}$ [62]. One can extend the above modularity conditions to the higher-genus case by considering the gapped boundary at generic Riemann surface, and studying the invariance of the corresponding boundary state under mapping class groups [61] [^7] Here, the boundary state is constructed by considering the TQFT on $\Sigma_{g}\times I$ where $\Sigma_{g}$ is the genus $g$ surface and $I$ is the interval, and then one end of the interval is filled with the gapped boundary. This defines a TQFT state on $\Sigma_{g}$ on the other end, which is the boundary state. Due to the topological nature of the gapped boundary, the boundary state has to be invariant under the mapping class group actions of $\Sigma_{g}$. However, such family of conditions within the genus <g<g still does not fully characterize the Lagrangian algebra. For instance, there are (1+1)D SPT phases with a certain finite group $G^{(g)}_{\text{sf}}$ which cannot be detected on the genus <g<g surface, but can be detected on genus $g$ (see Appendix B). This leads to a pair of gapped boundaries of the $G_{\text{sf}}^{(g)}$ gauge theory $(G_{\text{sf}}^{(g)},0),(G_{\text{sf}}^{(g)},\omega)$ that cannot be distinguished from the boundary state with genus <g<g.
@@ V.1 Application: Oblique SSB phase of $\mathrm{Rep}(G_{\mathrm{sf}})$ symmetry in (1+1)D
Nevertheless, these two phases with fully broken $\text{Rep}(G_{\text{sf}})$ are distinct phases. In particular, if one SSB phase is continuously connected to the other by a deformation preserving $\text{Rep}(G_{\text{sf}})$ symmetry, it has to go through a phase transition. This can be most easily seen by gauging the $\text{Rep}(G_{\text{sf}})$ symmetry [68], then the path connecting the SSB phases becomes the path from a trivial to nontrivial SPT phase preserving the dual $G_{\text{sf}}$ symmetry. This has to experience a phase transition. [^8] We thank Nat Tantivasadakarn for helpful discussions on this point. See [69] for explicit lattice models of distinct SSB phases of $\text{Rep}(G_{\text{sf}})$ symmetry.
@@ VI Higher dimensions: $Q_{8}$ gauge theory in (3+1)D
Since this slant product is trivial in cohomology $H^{2}(B{\mathbb{Z}}_{4},U(1))$, this $\nu=1$ phase does not induce permutation of the magnetic flux $\mathbf{i}$]. However, the above slant product leads to a subtle algebraic mixture among SPT defects, the magnetic fluxes and Wilson lines [50]. Suppose we have the junction of four SPT defects labeled by $\nu=1$ into $\nu=4$. If the magnetic surface operator labeled by the conjugacy class $\mathbf{i}]$ intersects with this junction, the intersection has to emit a ${\mathbb{Z}}_{4}$ electric Wilson line $\exp(\frac{2\pi i}{4}\int a)$ carrying the charge of ${\mathbb{Z}}_{4}=Z(\mathbf{i})$. Since only the 2d irreducible representation of $Q_{8}$ can carry the ${\mathbb{Z}}_{4}$ charge under $Z(\mathbf{i})$, one can say that this junction emits the 2d Wilson line of $Q_{8}$. This is a non-invertible categorical symmetry that involves the junction of 0-form SPT defects, magnetic fluxes and the 2d non-invertible Wilson line. See Fig. 9 for illustrations. The whole discussion about the $\nu=1$ defect above is symmetric under permuting group elements among $\mathbf{i},\mathbf{j},\mathbf{k}$.
@@
More generally, when $G$ is a discrete subgroup of $SU(2)$, one can always find a topological response $\omega\in Z^{3}(BG,U(1))$ which is nontrivial in cohomology $\omega]\neq 0$, but its slant product with any group element $g\in G$ becomes completely trivial $i_{g}\omega=0$ at the cochain level [71]. This implies that this class of finite groups has an analogue of soft symmetry, whose only nontrivial feature is an action on the point junctions of magnetic surfaces by a phase factor.
Figure 9: The junction of four $\nu=1$ defects into a $\nu=4$ defect. When the non-invertible magnetic surface operator $\mathbf{i}]=[\mathbf{j}]=[\mathbf{k}]$ intersects with this junction, their intersection emits an electric charge carried by the 2d Wilson line of $Q_{8}$ gauge theory.
@@ VII Discussion
Finally, it remains unknown if the generic soft symmetry has the $H^{3}$ obstruction characterized by $\mathcal{O}_{3}]=\mathcal{H}^{3}(\text{Aut}_{br}(\mathcal{C}),\mathcal{A})$. While the example of the soft symmetry provided in the main text does not have the $H^{3}$ obstruction, it remains unclear if the generic soft symmetries including the outer automorphism $\text{Out}_{\text{sf}}(G)$ carry the trivial $H^{3}$ obstruction or not. [^10] We note that [41] argued that the $H^{3}$ obstruction must be trivial when the symmetry does not permute anyons. Their argument assumes that the fusion multiplicities satisfy $N^{ab}_{c}\in\{0,1\}$ for all fusion vertices, and that the $F$-symbol $(F^{abc}_{d})_{ef}$ is always nonzero whenever the four relevant fusion spaces $V^{ab}_{e}$, $V^{ec}_{d}$, $V^{af}_{d}$, and $V^{bc}_{f}$ are all nonempty. We note, however, that there are modular tensor categories in which certain $F$-symbols vanish even though this is not forbidden by the fusion rules; for example, $F^{111}_{1}$ vanishes in $SU(2)_{4}$. Therefore, it is still unknown in general if $H^{3}$ obstruction is vanishing for soft symmetries. We thank Parsa Bonderson for bringing this point to our attention. See Refs. [3, 73] for examples of the outer automorphism in finite gauge theories with $H^{3}$ obstructions. This is an interesting question since it is often believed that the symmetry that does not permute anyons leads to trivial $H^{3}$ obstruction [3, 41].
@@ Appendix A Symmetric representative of second cohomology $H^{2}(B{G_{\mathrm{sf}}},U(1))$
Note that $\omega_{\text{sf}}$ gives another representative of $H^{2}(B{G_{\mathrm{sf}}},U(1))$ with $\omega_{\text{sf}}]=[\omega_{pq}]$. Let us explicitly construct such a function $\chi$. For convenience, a group element $g\in{G_{\mathrm{sf}}}$ is sometimes expressed as $g=g_{A}g_{B}$, where $g_{A}=p^{\alpha}q^{\beta}r_{1}^{\gamma_{1}}r_{2}^{\gamma_{2}}$ and $g_{B}=x^{\mu}y_{1}^{\nu_{1}}y_{2}^{\nu_{2}}$. We define a subset $\Gamma_{A}\subset{G_{\mathrm{sf}}}$ that consists of group elements $\Gamma_{A}=\{p^{\alpha}q^{\beta}r_{1}^{\gamma_{1}}r_{2}^{\gamma_{2}}\}$ with generic choices of $\alpha,\beta,\gamma_{1},\gamma_{2}$. Let us similarly define the subgroup $\Gamma_{B}\subset{G_{\mathrm{sf}}}$ where $\Gamma_{B}=\{x^{\mu}y_{1}^{\nu_{1}}y_{2}^{\nu_{2}}\}$.
@@
We consider the group $G^{(g)}_{\text{sf}}$ of order $2^{g^{2}+2g-1}$ with $g\geq 2$ which is the central extension ${\mathbb{Z}}_{2}^{g^{2}-1}\to G^{(g)}_{\text{sf}}\to{\mathbb{Z}}_{2}^{2g}$. The generators of ${\mathbb{Z}}_{2}^{2g}$ are labeled by $p_{j},q_{j}$ with $1\leq j\leq g$, and generators of ${\mathbb{Z}}_{2}^{g^{2}-1}$ are labeled by $x_{j}$ with $1\leq j\leq g-1$, and $y_{jk},z_{jk}$ with $1\leq j<k\leq g$. The central extension is characterized by
$\begin{split}[p_{1},q_{1}]&=x_{1},\quad[p_{g},q_{g}]=x_{g-1},\quad\\ [p_{j},q_{j}]&=x_{j-1}x_{j};\ 2\leq j\leq g-1\\ [p_{j},p_{k}]&=y_{jk};\ j<k\\ [p_{j},q_{k}]&=z_{jk};\ j<k\end{split}$ (67)
where $p,q]=pqp^{-1}q^{-1}$ is a group commutator. Note that the above group $G^{(g)}_{\text{sf}}$ with $g=2$ corresponds to the group $G_{\mathrm{sf}}$} discussed in the main text. Let us consider the SPT action $\omega^{(g)}\in Z^{2}(BG^{(g)}_{\text{sf}},U(1))$ characterized by
@@
where $n_{p_{1}}(h)\in Z^{1}(BG^{(g)}_{\text{sf}},U(1))$ is the mod 2 number of $p_{1}$ that appears in the expression of $h\in G^{(g)}_{\text{sf}}$ in terms of the product of group generators. Similar for $n_{q_{1}}(h)$. This SPT phase is detected on a genus $g$ surface with holonomy $(p_{j},q_{j})$, $1\leq j\leq g$ on each pair of cycles, while cannot be detected on closed oriented surfaces of genus <g<g.